OpenStudy (anonymous):

Solve the following systems of equations using the substitution method. What type of system is it? Name the solution if there is one. x=6y+2 3x-18y=4

OpenStudy (shinalcantara):

x=6y+2 ----equation 1 3x-18y=4 ----equation 2 substitute equation 1 to equation 2 3(6y+2)-18y=4 distribute 3 3(6y) +3(2)-18y=4 18y+6-18y=4 18y-18y=4-6 0 = -2 since it happened that the variable cancelled out therefore, the equations are equal making the solution set for both x and y infinite..

OpenStudy (shinalcantara):

oh i got it wrong.. since zero is not equal to -2 then there are no solutions making the system inconsistent